# Trapping of Single Atoms in Metasurface Optical Tweezer Arrays

### v2

@@ Abstract
Going beyond the example of a lens, we simulate optical tweezer arrays that can be generated with a pixel-based device (see Fig. 5c). For a smaller pixel size $d$, the effective NA attainable is higher, and thus the individual tweezers can be focused more tightly. This enhances the quality of the individual tweezers and allows the metasurfaces to accommodate more tweezer traps in the same area. In a series of simulations, we vary the pixel size $d$, but keep the device pixel count at $300\times 300$, while optimizing the phase mask to generate a $3\times 3$ square array of traps with 5 µm spacing in the focal plane (see Methods). As shown in Fig. 5c, the full-width-half-maximum (FWHM) of the traps steadily decreases as the pixel density $1/d$ increases. As a result, holographic metasurfaces with subwavelength pixels can generate optical tweezer arrays that are sufficiently tight at the focal plane for direct trapping of atoms, while devices with larger pixels require additional demagnification optics.
@@
Because of the small pixel sizes of holographic metasurfaces, a large number of pixels can be accommodated in a relatively small device footprint. This is advantageous as the number of high-quality traps that can be generated is positively correlated with the number of available pixels. To quantify this relation, we consider devices with pixel counts ranging from $000\times 1,$ 000 × 1 , 000 $(10^{6})$ to $000\times 16,$ 000 × 16 , 000 $(256\times 10^{6})$ and investigate the uniformity of trap intensity across the array as a function of the number of tweezer traps (see Methods). The results in Fig. 5d show that for a fixed pixel count, the uniformity drops monotonically when the required number of traps increases. The results also suggest a rule-of-thumb that $\sim 300$ pixels are needed to produce one high-quality tweezer trap in an array. For example, the number of highly uniform (uniformity $>95\%$) tweezer traps that can be generated by a top-end SLM with $4000\times 4000$ pixels will be fundamentally limited to $\sim 50,000$; the 2.32 mm metasurface in this work contains $\sim 8000\times 8000$ pixels, allowing the creation of $200,000$ highly uniform tweezer traps, provided that sufficient laser power is available.
@@ Conclusions
In this work, we have demonstrated single-atom trapping in metasurface optical tweezer arrays. We show that the uniformity of the arrays is comparable to that realized using existing methods; further improvements are expected via optimized fabrication and system integration. Because of subwavelength pixel sizes, holographic metasurfaces can reach a high effective NA. This allows for the creation of tightly focused tweezer arrays at the metasurface’s focal plane, enabling direct trapping of atoms without the need for demagnification or relay optics. Because of high pixel counts [44] and high power handling capabilities [36], metasurfaces offer a realistic path towards atomic tweezer arrays with $>100,000$ sites, overcoming fundamental limitations of existing approaches. This addresses a critical need for future applications in quantum simulation, quantum computing, quantum sensing, and optical clocks based on atomic arrays.
We envision future extensions of the metasurface optical tweezer array platform. By combining the static arrays demonstrated here with a dynamic sorting beam, it will become possible to rearrange atoms and create unity filled arrays. In addition, metasurfaces can be functionalized in various ways: resonant metasurfaces [54] can be designed to only impart a phase pattern on light within a narrow spectral band. This will enable the creation of metasurface stacks, comprised of several metasurfaces for different spectral bands, through which atomic arrays can be generated, sorted, and imaged. Additionally, metasurfaces can be designed to produce the exact same array pattern at different wavelengths, which may prove useful for both single- and dual-species atomic tweezer systems [11, 12]. Furthermore, polarization sensitive metasurfaces offer an additional tuning knob [32]. Finally, the passive stability and compactness of metasurfaces offers a robust platform for field-deployable atomic tweezer quantum devices.
@@ Acknowledgments
We are grateful to Xiaoyan Huang and Weijun Yuan for contributions in the early development of this project and Stephanie C. Malek for help with the development of the SRN material platform. We thank Chun-Wei Liu and Siwei Zhang for experimental assistance, and Dmytro Filin and Marianna Safronova for providing data on the optical polarizability of Sr atoms. We thank Ana Asenjo-Garcia, Stuart Masson, and Ricardo Gutierrez-Jauregui for fruitful discussions, and Tarik Yefsah for critical reading of the manuscript. This work was supported by the National Science Foundation (Award nos. 1936359, 2040702, and 2004685) and the Air Force Office of Scientific Research (Award nos. FA9550-16-1-0322 and FA9550-23-1-0404). Device fabrication was carried out at the Columbia Nano Initiative cleanroom, at the Advanced Science Research Center Nanofabrication Facility at the Graduate Center of the City University of New York, and at the Center for Functional Nanomaterials, Brookhaven National Laboratory, supported by the US Department of Energy, Office of Basic Energy Sciences (Contract no. DESC0012704). B.S. acknowledges support from the National Research Foundation of Korea (Award no. 2021M3H3A1036573). S.W. acknowledges support from the Alfred P. Sloan Foundation.
@@
All authors contributed substantially to the work presented in this paper. A.H., X.S., M.W., and B.S. carried out the atomic experiments. Y.X. and J.W. designed and fabricated the metasurfaces. N.Y. and S.W. supervised the study. All authors contributed to the data analysis and writing of the paper.
@@ Methods
Material platform for visible-spectrum metasurfaces
To fabricate holographic metasurfaces operating at a wavelength of $\lambda=520$ nm, titanium dioxide ($\text{TiO}_{2}$) is frequently selected due to its high refractive index and minimal absorption [55, 56]. Various methods have been explored for creating $\text{TiO}_{2}$-based metasurfaces, such as atomic layer deposition (ALD) on patterned electron-beam (e-beam) resist [55] or employing inductively coupled plasma (ICP) etching with a patterned chromium (Cr) mask [32]. However, these methodologies face significant scalability and integration challenges within complementary metal-oxide-semiconductor (CMOS) workflows due to the slow deposition rates of ALD or the ICP etching of $\text{TiO}_{2}$ and eventual removal of Cr masks.
Considering these challenges, we employ a novel material platform, silicon-rich silicon nitride (SRN), to fabricate our metasurfaces. SRN has been used to build nonlocal metasurfaces with great performance and excellent wavelength selectivity, enabled by photonic resonances in the visible spectrum [57]. We utilize plasma-enhanced chemical vapor deposition (PECVD) to deposit SRN films onto optically thick fused silica substrates. This enables precise control over the real part of the linear refractive index of SRN films, which can be adjusted from 1.92 to 3.1 by varying the ratios of the precursor gases ($\text{SiH}_{4}$, $\text{N}_{2}$ and $\text{NH}_{3}$) used in the PECVD process [58]. Such tunability provides the flexibility required to accommodate a wide range of functionalities. For our specific application at $\lambda=520$ nm, we deposit SRN films with a real part of the refractive index of 2.3 and an imaginary part maintained near zero. This greatly enhances the design possibilities for these metasurfaces while ensuring high transmission and high diffraction efficiency.
@@
Extended Data Fig. 1: A calculated phase-only hologram and its corresponding focal plane trap pattern. a, A phase-only hologram generated by the modified Gerchberg-Saxton algorithm with $000\times 4,$ 000 × 4 , 000 pixel resolution and a 290 nm pixel size. b, Simulated intensity distribution at the focal plane, which shows the Statue of Liberty made of 183 traps as designed.
Calculation and optimization of phase-only holograms
@@
To compute the phase-only holograms used to create arbitrary trap arrays with a designated focal length and NA, we employ a Gerchberg-Saxton algorithm-based approach [59]. The original algorithm utilizes the Rayleigh-Sommerfeld diffraction integral to iterate between the diffractive optical elements (DOE) plane (in our case, the metasurface plane) and the focal plane by forward and backward light propagation. This iterative process ultimately determines the phase profile necessary to generate the desired pattern. In each iteration, the input amplitude at the metasurface plane is set to unity, enforcing a phase-only condition for the algorithm. An example of the resulting phase-only hologram and the corresponding simulated trap array pattern is shown in Extended Data Fig. 1.
@@
Extended Data Fig. 2: RCWA-simulated phase response and transmission of the selected meta-unit library. a, Phase response of the meta-unit library as a function of pillar width $a$. b, Transmission of the meta-unit library. All individual meta-units have a transmission over $95\%$ (marked as the grey dashed line).
@@
However, the standard Gerchberg-Saxton algorithm is designed for general holography and inevitably introduces speckle and noise into the resulting patterns. Although this can be mitigated by using soft operations when updating the hologram [60], achieving high-quality, point-like traps remains a significant challenge. To address this, we employ a weighted Gerchberg-Saxton (GSW) algorithm [61], modified to account for the subwavelength pixel size and the direct focusing capability of our meta-holograms. At the $(k+1)^{\text{th}}$ iteration, the target amplitude distribution is determined by the simulated focal plane amplitude and the target amplitude from the previous iteration via
$A^{k+1}_{\mathrm{target}}(x,y)=\frac{A^{k}_{\mathrm{target}}(x,y)}{A^{k}_{\mathrm{simulated}}(x,y)},$
@@
where $A^{k+1}_{\mathrm{target}}(x,y)$ and $A^{k}_{\mathrm{target}}(x,y)$ represent the target amplitude distributions at the $(k+1)^{\text{th}}$ and $k^{\text{th}}$ iterations, respectively, and $A^{k}_{\mathrm{simulated}}$ is the simulated focal plane amplitude. To prevent the target amplitude from producing extreme numerical values, numerical normalization and soft operations are performed in each update. Through this approach, we can generate high-quality phase-only holograms capable of producing near-uniform point trap arrays in arbitrary geometries, with a simulated standard deviation below $3\%$ even for arrays consisting of hundreds of traps.
Meta-unit library design
To construct the meta-unit library for our metasurfaces, we utilize rigorous coupled-wave analysis (RCWA) simulations to determine the phase and amplitude response of individual SRN nanopillars. By using non-birefringent meta-units and constraining the meta-units to square-shaped cross-sections, we maximize the metasurfaces’ diffraction efficiency while minimizing fabrication defects.
@@
The pitch between the nanopillars (pixel size) is set to 290 nm. The height of the nanopillars is set to 750 nm and their width ranges from 100 to 190 nm, which provides comprehensive phase coverage over the $2\pi$ range while maintaining near unity transmission. The subwavelength size enables an achievable numerical aperture (NA) of up to 0.9, allowing great flexibility in designing the phase hologram. Extended Data Fig. 2 shows the RCWA-simulated phase and amplitude response for the selected meta-unit library, which consists of 14 SRN nanopillars of varying square cross-sectional sizes, all with a uniform height of 750 nm. Due to the subwavelength pixel size, this small library is sufficient to achieve fine-grained phase control across a $\lambda^{2}$ area. To prevent defects caused by missing or collapsed pillars during fabrication, the smallest pillar is designed with a minimum edge width of 100 nm. The library achieves a high overall transmission that exceeds $95\%$ across the entire set of meta-units, resulting in an experimental diffraction efficiency of over $60\%$.
@@
Fabrication of SRN metasurfaces
@@
The fabricated metasurfaces feature circular shapes with diameters ranging from 1.16 mm ($000\times 4,$ 000 × 4 , 000 pixel resolution with a 290 nm size) to 2.32 mm ($000\times 8,$ 000 × 8 , 000 pixel resolution with a 290 nm pixel size). As shown in Extended Data Fig. 3b, the scanning electron microscope (SEM) images of the fabricated samples reveal high-quality SRN nanopillar matrices with minimal defects and clean side walls.
@@
The SRN metasurfaces manufactured in this work are large enough to be easily visible by eye (Extended Data Fig. 4). We choose a circular footprint to facilitate alignment with the incident beam. This also reduces unwanted diffraction off the edges of the metasurface.
@@
Metasurfaces have a high forward scattering efficiency, enabling more tweezers to be generated given a fixed input power. The portion of power from the input beam diffracted by the metasurface depends on the input beam shape and the efficiency of the metasurface itself. We choose a beam size slightly larger than the metasurface to ensure a flat intensity profile, which our metasurface design is based on. Therefore, some input power transmits through the substrate without being modulated. We choose an input beam diameter of 1.5 mm for the metasurfaces with 1.16 mm diameter. Excluding the non-overlapped portion of the beam, we observe a diffraction efficiency of $60\%$ into the first order. In the future, the phase mask can be designed for a Gaussian input, allowing the full incident power to be used.
@@
The metasurface platform also has a high power handling capability. We have tested the damage threshold by maximizing the input power and decreasing the waist of the incident beam. Above the threshold, the largest meta-atoms (190 nm) show signs of thermal expansion, closing the air gap that exists between neighboring pillars. We observe the current intensity handling capabilities of the SRN metasurfaces to be up to $25$ W/mm2, about 5×5$\times$5 × larger than the damage threshold of liquid crystal SLMs. By reducing the average pillar size it should be possible to improve this threshold further. Also, alternative materials promise even better power handling. For example, we have observed TiO2 metasurfaces to display no signs of degradation up to 36 W/mm2.
@@
Given the measured SRN damage threshold, we provide an estimate for how many traps can realistically be generated with the current parameters. Assuming a 2.3 mm $\times$ 2.3 mm device, 1 mW per trap, and $60\%$ diffraction efficiency, we can in principle generate approximately $80,000$ tweezers; therefore, metasurface tweezer arrays with $>100,000$ sites are realistic without major improvements over our current technology. With modest improvements in the power handling and for device footprints of 10 mm $\times$ 10 mm, arrays with $>1,000,000$ traps are within realistic reach.
@@
Characterizing the uniformity of the metasurface array by measuring the peak optical intensity of each trap, we find a non-uniformity of $4\%$. This is measured by collimating the metasurface pattern with a single high NA microscope lens, without the full relay optics used in this work. This is notably better than the non-uniformity of $7.5\%$ measured with the atoms. We attribute this discrepancy to imperfections in the relay optics paths. Removal or improvement of the relay optics should allow a $4\%$ non-uniformity of the atomic traps. Of note, we were able to rapidly improve the intensity non-uniformity from an initial 20% to single digit percentages via improvements in the design and fabrication. This uniformity is insensitive to slight misalignment of the incident angle into the metasurface and independent of the pattern geometry.
@@
The preparation of ultracold strontium atoms starts with the generation of a cold atomic beam with a 2D magneto-optical trap (MOT) that is directly loaded from resistively heated dispensers, as described in previous work [62]. The atoms are transferred to a glass cell vacuum chamber via a push-beam where they are captured and cooled to about 1 mK by a 3D MOT operating on the 461 nm transition. We simultaneously operate two repump lasers at 679 nm and 707 nm to close a loss channel present in the 461 nm cooling scheme. To further cool the atoms, we then operate a second MOT on the narrow-line 689 nm transition. We begin the transfer between MOTs by frequency broadening the 689 nm line to 3 MHz to match the temperature distribution of the 461 nm MOT, before smoothly narrowing down to a single-frequency 689 nm MOT with on average $10^{5}$ atoms at 1 µK.

### v3

@@ Abstract
Going beyond the example of a lens, we simulate optical tweezer arrays that can be generated with a pixel-based device (see Fig. 5c). For a smaller pixel size $d$, the effective NA attainable is higher, and thus the individual tweezers can be focused more tightly. This allows the metasurfaces to accommodate more tweezer traps in the same area. In a series of simulations, we vary the pixel size $d$, but keep the device pixel count at $300\times 300$, while optimizing the phase mask to generate a $3\times 3$ square array with 5 µm spacing in the focal plane (see Methods). As shown in Fig. 5c, the full-width-half-maximum (FWHM) of the traps steadily decreases as the pixel density $1/d$ increases. As a result, holographic metasurfaces with subwavelength pixels can generate optical tweezer arrays that are sufficiently tight at the focal plane for direct trapping of atoms, while devices with larger pixels require additional demagnification optics.
@@
The small pixel size of metasurfaces also allows a large number of pixels to be accommodated within a compact device footprint. This is advantageous as the number of high-quality traps that can be generated is positively correlated with the number of pixels. To quantify this relation, we consider devices with pixel counts ranging from $1,000\times 1,000$ $(10^{6})$ to $16,000\times 16,000$ $(256\times 10^{6})$ and investigate the uniformity of trap intensity across the array as a function of the number of tweezer traps (see Methods). The results in Fig. 5d show that for a fixed pixel count, the uniformity drops monotonically when the number of traps increases. The data suggests a rule-of-thumb that $\sim 300$ pixels are needed to produce one high-quality tweezer trap in an array  [^1] It is important to note this is a fundamental limitation that cannot be improved upon, for example, via active feedback in SLMs or DMDs.. For example, the number of highly uniform (uniformity $95\%$) tweezer traps that can be generated by a top-end SLM with $4000\times 4000$ pixels will be fundamentally limited to $\sim 50,000$; in contrast, the pixel count of metasurfaces can be well beyond $8000\times 8000$, allowing the creation of arrays with >200,000>200,000 traps, provided that sufficient laser power is available.
360,000 Tweezer Array
As a demonstration of the high scalability of metasurface arrays, we experimentally realize an array with 360,000 tweezer traps (see Fig. 6). The traps are arranged on a $600\times 600$ square lattice with a spacing of 2.5 µm between nearest-neighbor sites. The metasurface has a diameter of $3.5$ mm, contains approximately 114 million pixels, and is made of titanium dioxide, offering exceptional power-handling capabilities (see Methods).
We record a high-resolution image of the full array and characterize the uniformity of the tweezer intensities. Using the setup shown in Fig. 6a, which has a high-resolution field of view of about 190 x 140 µm, we raster scan the array, which has a real space size of 1.5 mm x 1.5 mm, and stich together a composite image from more than 100 individual images. The full image frame, with a $50\times$ 4K resolution, is provided in the SI. Exemplary high-resolution images of the edges and corners of the array are shown in Fig. 6b, demonstrating the high quality and uniformity of the tweezers across the array. From a quantitative analysis of the full array, we find a uniformity of the tweezer power of $92\%$ (Fig. 6c). With a moderate increase of metasurface size, arrays with >>1,000,000 traps are within realistic reach.
@@ Conclusions
In this work, we have demonstrated single-atom trapping in metasurface optical tweezer arrays. We show that the uniformity of the arrays is comparable to that realized using existing methods; additional improvements are expected by optimizing fabrication and system integration. Because of subwavelength pixel sizes, holographic metasurfaces can reach a high effective NA. This allows for the creation of tightly focused tweezer arrays at the metasurface’s focal plane, enabling direct trapping of atoms without demagnification or relay optics. With their high pixel counts [46] and outstanding power-handling capabilities [37], metasurfaces offer a realistic path towards tweezer arrays with more than $100,000$ atoms. This addresses a critical need for future applications in quantum simulation, quantum computing, quantum sensing, and optical clocks based on atomic arrays.
We envision future extensions of the metasurface optical tweezer array platform. By combining the static arrays demonstrated here with a dynamic sorting beam, it will become possible to rearrange atoms and create unity filled arrays. In addition, metasurfaces can be functionalized in various ways: resonant metasurfaces [55] can be designed to only impart a phase pattern on light within a narrow spectral band, while leaving wavelengths outside of this band unchanged. This will enable the creation of multifunctional metasurfaces through which atomic arrays can be generated, sorted, and imaged. In addition, by employing wavelength or polarization multiplexing  [33], metasurfaces can be designed to produce the same tweezer array pattern at different wavelengths, which may prove useful for both single- and dual-species atomic systems  [11, 12]. Finally, the development of active metasurfaces that allow for feedback and rearrangement of atomic arrays constitute an intriguing frontier for future research [56].
@@ Acknowledgments
We are grateful to Xiaoyan Huang and Weijun Yuan for contributions in the early development of this project and Stephanie C. Malek for help with the development of the SRN material platform. We thank Chun-Wei Liu and Siwei Zhang for experimental assistance, and Dmytro Filin and Marianna Safronova for providing data on the optical polarizability of Sr atoms. We thank Ana Asenjo-Garcia, Stuart Masson, and Ricardo Gutierrez-Jauregui for fruitful discussions, and Tarik Yefsah for critical reading of the manuscript. This work was supported by the National Science Foundation (Award nos. 1936359, 2040702, and 2004685) and the Air Force Office of Scientific Research (Award nos. FA9550-16-1-0322 and FA9550-23-1-0404). Device fabrication was carried out at the Columbia Nano Initiative cleanroom, at the Advanced Science Research Center Nanofabrication Facility at the Graduate Center of the City University of New York, and at the Center for Functional Nanomaterials, Brookhaven National Laboratory, supported by the US Department of Energy, Office of Basic Energy Sciences (Contract no. DESC0012704). B.S. acknowledges support from the National Research Foundation of Korea (Award no. 2021M3H3A1036573). S.W. acknowledges support from the Alfred P. Sloan Foundation. N.Y.  acknowledges the Gordon and Betty Moore Foundation Experimental Physics Investigators (EPI) initiative (Award no. 11561).
@@
All authors contributed substantially to the work presented in this paper. A.H., X.S., M.W., and B.S. carried out the atomic experiments. Y.X., J.W., and Z.Z. designed and fabricated the metasurfaces. N.Y. and S.W. supervised the study. All authors contributed to the data analysis and writing of the paper.
@@ Methods
Material platforms for visible-spectrum metasurfaces
For the use of metasurfaces in atom trapping, it is beneficial to have several photonic material platforms available to accommodate different laser wavelength requirements, power-handling capabilities, and available fabrication workflows. We implement and benchmark two CMOS-compatible materials platforms, both supporting the visible spectrum:
Titanium-dioxide ($\text{TiO}_{2}$): Titanium-dioxide metasurfaces leverage $\text{TiO}_{2}$’s high refractive index ($n\gtrsim 2.4$ at a wavelength of 520 nm) and negligible absorption to achieve high diffraction efficiencies and low losses across a broad band of wavelengths  [39, 57, 58]. Established fabrication approaches include (i) conformal atomic-layer deposition (ALD) of $\text{TiO}_{2}$ on patterned electron-beam resist [39], and (ii) inductively-coupled-plasma (ICP) etching structures in a sputtered $\text{TiO}_{2}$ thin film using a Cr hard mask [33]. While these methods deliver outstanding optical performance, their throughput is limited by the slow ALD growth and by challenges associated with the hard-mask liftoff, motivating a different platform with faster fabrication for large-scale production.
Silicon-rich $\text{Si}_{3}\text{N}_{4}$ (SRN): We introduce silicon-rich silicon nitride (SRN) as a new metasurface platform [59]. SRN thin films are deposited by high-rate plasma-enhanced chemical-vapour deposition (PECVD) directly onto fused-silica wafers. By varying the ratios of the precursor gases ($\text{SiH}_{4}$, $\text{N}_{2}$, and $\text{NH}_{3}$) used in the PECVD process [60], we can precisely control the real part of the complex refractive index of the SRN films, which can be adjusted from 1.9 to 3.1. This tunability enables versatile designs while retaining fully CMOS-compatible blanket deposition and etch steps - attributes that have already been exploited for the fabrication of large-aperture metalenses with $\text{NA}\approx 0.98$ in the visible regime [61]. In this work, we use 750-nm-thick SRN films with $n=2.3$ and negligible extinction. As detailed below, the resulting SRN metasurfaces achieve high forward-scattering efficiency and withstand high continuous intensities of at least 25 W/mm2. For applications requiring even higher power handling, $\text{TiO}_{2}$ offers extreme robustness, tolerating intensities above 2,000 W/mm2, which is especially beneficial for scaling of tweezer arrays beyond 100,000 traps.
@@
Extended Data Fig. 1: A calculated phase-only hologram and its corresponding focal-plane trap pattern. a, A phase-only hologram generated by the modified Gerchberg-Saxton algorithm with $4,000\times 4,000$ pixels and a pixel size of 290 nm. b, Simulated intensity distribution at the focal plane, showing a Statue of Liberty pattern consisting of 183 traps.
Design and optimization of phase-only holograms
@@
To design phase-only holograms that create tweezer arrays with a specific focal length and NA, we use an approach based on the Gerchberg-Saxton algorithm [62]. The original algorithm utilizes the Rayleigh-Sommerfeld diffraction integral to iterate between the device plane (in our case, the metasurface plane) and the focal plane by forward and backward light propagation. This iterative process converges to a phase profile that produces the desired focal-plane intensity distribution. In each iteration, the input amplitude over the metasurface plane is set to unity, enforcing a phase-only condition for the algorithm.
@@
Extended Data Fig. 2: RCWA-calculated phase response and transmission of the SRN meta-atom library consisting of 13 nanopillars with different cross-sectional sizes. a, Phase response of the meta-atom library as a function of nanopillar width $a$. b, Transmission or forward scattering efficiency of the meta-atom library. All meta-atoms have a transmission over $95\%$ (marked as the grey dashed line).
@@
The standard Gerchberg-Saxton algorithm is designed for general holography and inevitably introduces speckle and noise into the resulting optical patterns. Although this can be mitigated by using "soft operations" when updating the hologram [63], achieving high-quality, point-like traps remains a significant challenge. To address this, we employ a weighted Gerchberg-Saxton (GSW) algorithm [64], modified to account for the subwavelength pixel size and the direct focusing capability of our metasurface holograms. At the $(k+1)^{\text{th}}$ iteration, the target amplitude distribution is determined by the calculated focal plane amplitude and the target amplitude from the previous iteration via
$A^{k+1}_{\mathrm{target}}(x,y)=\frac{A^{k}_{\mathrm{target}}(x,y)}{A^{k}_{\mathrm{calculated}}(x,y)},$
@@
where $A^{k+1}_{\mathrm{target}}(x,y)$ and $A^{k}_{\mathrm{target}}(x,y)$ represent the target amplitude distributions at the $(k+1)^{\text{th}}$ and $k^{\text{th}}$ iterations, respectively, and $A^{k}_{\mathrm{calculated}}$ is the calculated focal plane amplitude. To prevent the target amplitude from producing extreme numerical values, numerical normalization and "soft operations" are performed in each update. Through this approach, we can generate high-quality phase-only holograms capable of producing near-uniform point trap arrays in arbitrary geometries, with a simulated standard deviation below $3\%$ even for arrays consisting of >100,000>100,000 traps. An example of a phase-only hologram and the corresponding simulated tweezer array is shown in Extended Data Fig. 1.
Design of Meta-atom library
@@
To construct the meta-atom library for our metasurfaces, we utilize rigorous coupled-wave analysis (RCWA) to calculate the phase and amplitude response of individual SRN or $\text{TiO}_{2}$ nanopillars. The meta-atoms are non-birefringent and have square-shaped cross-sections. The center-to-center distance between adjacent nanopillars (i.e., the pixel size of the meta-atoms) is set to 290 nm. The height of the nanopillars is set to 750 nm for SRN and 600 nm for $\text{TiO}_{2}$ and their width ranges from 100 nm to 190 nm to provide a comprehensive phase coverage over the $2\pi$ range at a wavelength of 520 nm, while maintaining near-unity transmission or forward scattering efficiency (Extended Data Fig. 2). The shown library achieves a high overall transmission that exceeds 95% across the entire set of meta-units, resulting in an experimental diffraction efficiency of over 60%. To prevent defects caused by missing or collapsed pillars during fabrication, the smallest nanopillar has an edge width of 100 nm. The subwavelength width of the nanopillars and size of the meta-atoms allow for high-resolution sampling of the designed metasurface phase profile with an achievable NA of 0.9.
@@
Fabrication of $\text{TiO}_{2}$ and SRN metasurfaces
Fabrication of $\text{TiO}_{2}$ metasurfaces: Metasurfaces made from ALD-grown TiO2 have been demonstrated as a promising platform for diffraction-limited focusing [39], broadband phase control [65, 58], and complex beam shaping in the visible [66, 67, 68]. Our $\text{TiO}_{2}$ metasurfaces are fabricated on 0.5-mm-thick, double-side-polished fused-silica wafers. A 700-nm ZEP-520A layer is spin-coated and baked (180C180\ $^{\circ}\text${C}, 3 min). The thickness of the resist is verified with a stylus profiler (KLA P-17). After applying an E-spacer layer, the nano-pillar template is written by 100 kV electron-beam lithography (EBL; Elionix ELS-G100) with a current of 2 nA and a step size of 4 nm. The resist is developed in chilled o-Xylene, rinsed in IPA, and nitrogen-dried, yielding apertures whose depth sets the final $\text{TiO}_{2}$ pillar height. Amorphous $\text{TiO}_{2}$ is then conformally deposited at 200C200\ $^{\circ}\text${C} in an ALD reactor (Cambridge NanoTech Savannah) until the apertures are fully filled. Excess $\text{TiO}_{2}$ material on top is removed by inductively coupled plasma (ICP) etching (CHF3/Ar/O2, Oxford PlasmaPro 100 Cobra) down to the resist surface. A final downstream plasma ashing at 220C220\ $^{\circ}\text${C} (Matrix Plasma Asher) removes the resist template, leaving free-standing $\text{TiO}_{2}$ nanopillars on the fused-silica substrate.
@@
The fabricated $\text{TiO}_{2}$ and SRN metasurfaces have circular footprints with diameters ranging from 1.16 mm (approximately $4,000\times 4,000$ nanopillars or pixels) to 3.48 mm (approximately $12,000\times 12,000$ pixels; Extended Data Fig. 4a). Scanning electron microscope (SEM) images of the fabricated metasurfaces (Extended Data Fig. 3b) show that the nanopillars are defect-free and have vertical side walls.
@@
The metasurfaces have a circular footprint to facilitate their alignment with the incident beam and to reduce diffraction at the metasurface edges. The portion of the laser beam captured and shaped by the metasurface depends on the input beam size and the efficiency of the metasurface. We choose a beam size slightly larger than the metasurface to ensure a relatively uniform intensity profile and flat optical wavefront across the metasurface, which our metasurface design is based on. For example, the input beam diameter is approximately 1.5 mm for the metasurfaces with a diameter of 1.16 mm. Therefore, some input power transmits through the substrate without being modulated. Excluding the non-overlapping portion of the beam, we observe a diffraction efficiency of $60\%$ for our metasurfaces. In the future, the phase mask can be designed for a Gaussian input beam, allowing an even larger portion of incident power to be used.
The high forward scattering efficiency of metasurfaces enables a large number of tweezers for a given incident laser power. Unlike SLMs or AODs, where the 0th-order diffraction is not modulated and manifests as a bright specular spot, the non-modulated laser power from a metasurface is diffused over the full $4\pi$ steradians through subwavelength scattering, leaving the atomic plane essentially background-free. Underlining the high potential of our approach, recent demonstrations of dispersion-engineered and inverse-designed metasurfaces have already demonstrated an efficiency above $90\%$ over a wavelength range of hundreds of nanometers [58, 69], and a recent quasi-3D metasurface even demonstrated $99\%$ anomalous-refraction efficiency [70].
@@
The metasurface platform also has a high power-handling capability. We have tested the damage threshold by maximizing the input power and decreasing the waist of the incident beam. We observe that the damage threshold of the SRN metasurfaces is at least $25$ W/mm2, about $5\times$ larger than that of liquid-crystal SLMs. By reducing the "silicon richness" in SRN and thus decreasing the extinction coefficient of the material, it should be possible to further improve the power-handling capability of the SRN metasurfaces. Furthermore, alternative materials promise even better power handling; for TiO2 metasurfaces, we observe no signs of degradation up to 2000 W/mm2. A comparison of the properties of metasurfaces with AODs and SLMs is provided in Extended Data Table 1.
@@
Given the measured damage threshold of the SRN metasurfaces, we provide an estimate of how many traps can realistically be generated with the current device parameters. Assuming a 2.3 mm $\times$ 2.3 mm device with $60\%$ diffraction efficiency, a laser power of 130 W, and about 1 mW laser power per tweezer trap, approximately $80,000$ tweezers can be generated; therefore, metasurface tweezer arrays with >100,000>100,000 sites are realistic without major improvements over current technology. With a modest improvement on the power-handling capability and an increase in the device footprint to 10 mm $\times$ 10 mm, arrays with >1,000,000>1,000,000 traps are within realistic reach even with SRN. The power handling capabilities of TiO2 already far exceed the requirements to generate >1,000,000>1,000,000 traps.
@@
In the main text, we report the non-uniformity of the trap depth as $7.5\%$, measured with atoms. It is worth noting that the non-uniformity of the optical intensity of the traps, measured in the image plane of the metasurface, is significantly lower. By directly measuring the peak optical intensity of each tweezer spot in the image plane behind the metasurface (i.e., without the relay optics), we find a non-uniformity of $4\%$. We attribute the higher non-uniformity in the trap depth of the atomic array to imperfections that are introduced by the relay optics. Removal or improvement of the relay optics should allow for a $4\%$ non-uniformity of the atomic traps. Remarkably, we were able to rapidly improve the intensity non-uniformity from an initial $20\%$ to single digit percentages via improvements in the metasurface design and fabrication, and further potential for improvements is highly likely. The uniformity is insensitive to slight misalignment of the angle or position of the incident beam with respect to the metasurface and is independent of the specific array geometry.
Characteristics AOD SLM Metasurface
Type Active Active Passive
Relay optics Required Required Not required
Wavelength range Specified by model Specified by model
Widely tunable
Power handling $\sim 10$ W/mm2 $\sim 2$ W/mm2
>25>25 W/mm2 - SRN
>2000>2000 W/mm2 - TiO2
Diffraction efficiency $\sim 50\%$ $\sim 40\%$ $\sim 60\%$
Device footprint 1101-10 cm2 1101-10 cm2 1101-10 mm2
Pixel size NA 2102-10 µm 100500100-500 nm
Trap geometry
2D simple
geometry
Arbitrary pattern
in any dimension
Arbitrary pattern
in any dimension
Extended Data Table 1: Comparison of beam-shaping devices.
@@
The preparation of ultracold strontium atoms starts with the generation of a cold atomic beam with a 2D magneto-optical trap (MOT) that is directly loaded from resistively heated dispensers, as described in previous work [71]. The atoms are transferred to a glass cell vacuum chamber via a push-beam where they are captured and cooled to about 1 mK by a 3D MOT operating on the 461 nm transition. We simultaneously operate two repump lasers at 679 nm and 707 nm to close a loss channel present in the 461-nm cooling scheme. To further cool the atoms, we operate a second MOT on the narrow-line 689-nm transition. We begin the transfer between MOTs by frequency broadening the 689-nm line to 3 MHz to match the temperature distribution of the 461-nm MOT, before smoothly narrowing down to a single-frequency 689-nm MOT with on average $10^{5}$ atoms at 1 µK.